OpenStudy (anonymous):

factor the trinomial 3x^2-17x-6

OpenStudy (anonymous):

twilight i go it!

OpenStudy (anonymous):

3x^2-17x-6 -----> (3x ) ( x ) = 3x^2

OpenStudy (anonymous):

(3x+1) (x-6)

OpenStudy (anonymous):

times that stuff all together 3x*x= 3x^2 - 18x+1x-6 solve and get orignal

OpenStudy (anonymous):

https://www.youtube.com/watch?v=AMEau9OE6Bs

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

:D

OpenStudy (jhannybean):

\[3x^2-17x-6\]\[x^2 -17x -18\]\[(x-18)(x+1)\]\[\left(x-\frac{18}{3}\right)\left(x+\frac{1}{3}\right)\]Simplify whichever fractions can simplify, and then multiply the denominators of the fractions that do not simplify that easily. \[(x-6)(3x+1)\]

OpenStudy (jhannybean):

We divided our factored form by 3 because we initially multiplied it to the constant to make it easier to simplify.

OpenStudy (anonymous):

thank you @Jhannybean !

OpenStudy (jhannybean):

No problem :)